OpenStudy (anonymous):

Calculus problem: find the values of constants a, b, and c so that the graph y=(x^2+a)/(bx+c) has a local max at x=3 and a local min at (-1,2)

myininaya (myininaya):

Have you differentiated y yet?

OpenStudy (anonymous):

So far, I have...y=(9+a)/(3b+c) by subing in x=3 2b-2c-a=1 by subing in (-1,-2) and subbing in x=3 b=ab+2c from the first derivative and subbing in (-1,-2)

OpenStudy (anonymous):

2b-2c-a=1 by subing in (-1,-2) **

OpenStudy (anonymous):

c=(ab-9b)/6 from the first derivative and subbing in x=3

myininaya (myininaya):

we know we will have f'(x)=0 when x=3 and f'(x)=0 when x=-1 You are also given (-1,2) is on the graph of y=f(x) Also I'm slightly confused it the point (-1,2) or (-1,-2)?

OpenStudy (anonymous):

oops im sorry the point is (-1,-2)

myininaya (myininaya):

what did you get for y'?

OpenStudy (anonymous):

and yeah, i used that info and got equations... but the problem is i have to many unknowns so if i substitute the equations in, ill still have unknowns lol. y'=[x(bx+2c)-ab] /(bx+c)^2

myininaya (myininaya):

you will have 3 equations with 3 variables

myininaya (myininaya):

\[y'=\frac{(x^2+a)'(bx+c)-(bx+c)'(x^2+a)}{(bx+c)^2} =\frac{2x(bx+c)-b(x^2+a)}{(bx+c)^2} \\ y'=\frac{2bx^2+2cx-bx^2-ba}{(bx+c)^2}=\frac{bx^2+2cx-ba}{(bx+c)^2}\] your y' looks wonderful :)

myininaya (myininaya):

so we will have one equation given be f(-1)=-2 and two equations given by f'(x)=0 when x=3 then also at x=-1

myininaya (myininaya):

\[-2=\frac{1+a}{-b+c} \text{ from the } f(-1)=-2 \text{ thing } \\ 0=\frac{9b+6c-ba}{(3b+c)^2} \text{ from the } f'(3)=0 \\ 0=\frac{b-2c-ba}{(-b+c)^2} \text{ from the } f'(-1)=0\]

myininaya (myininaya):

those second two equations we can't assume the bottom doesn't equal in zero and rewrite the equations

myininaya (myininaya):

and let's get rid of the fraction for the first one too

myininaya (myininaya):

\[2b-2c=1+a \\ 0=9b+6c-ba \\ 0=b-2c-ba\]

myininaya (myininaya):

did you get this far?

OpenStudy (anonymous):

yup!

myininaya (myininaya):

\[0=-2b+2c+1+a \\ 0=9b+6c-ba \\ 0=b-2c-ba \] Well right away I can see how we can get a system of two equations with two variables

myininaya (myininaya):

\[0=9b+6c-ba \\ 0=b-2c-ba \\ \text{ multiply bottom equation by } -1 \\ 0=9b+6c-ba \\ 0=-b+2c+ba \\ \text{ add equations together } \\ 0=8b+4c \\ 8b=-4c \\ c=-2b \\ \text{ and the first equation was } 0=-2b+2c+1+a \\ \text{ so we can sub in } c=-2b \text{ in that one } \\ 0=-2b+2(-2b)+1+a\] and lied about the system of two equation with 2 variables forgot the first 1 had a in it

myininaya (myininaya):

but anyways this system should be solvable since we do have 3 equations with 3 unknowns we just have to play around with it

OpenStudy (anonymous):

alright! but wouldn't c=-b ?

myininaya (myininaya):

write the equation should be 0=8b+8c 8b=-8c b=-c

myininaya (myininaya):

so that equation should say 0=-2b+2(-b)+1+a

myininaya (myininaya):

0=-2b+1+a

myininaya (myininaya):

2b-1=a

myininaya (myininaya):

so you 2b-1=a and c=-b to write another equation in terms of just b

myininaya (myininaya):

that will allow us to solve for b

myininaya (myininaya):

you should get two possible values for b

myininaya (myininaya):

have you gotten those yet

OpenStudy (anonymous):

uhm im bit confused where to solve for b

myininaya (myininaya):

use the equations given to you for example you have 0=9b+6c-ba

myininaya (myininaya):

replace a with (2b-1) replace c with -b

myininaya (myininaya):

\[\text{ solve } 0=9b+6(-b)-b(2b-1) \\ 0=9b-6b-2b^2+b \\ 0=4b-2b^2 \]

myininaya (myininaya):

you will get two difference solutions for (a,b,c) but we will see later only one solution will work of the two

OpenStudy (anonymous):

i got b=0 and b=-1/2

myininaya (myininaya):

b=-1/2?

myininaya (myininaya):

let me see

OpenStudy (anonymous):

i think there was a mistake when subbing in b=-c into 2b-2c-a-1=0, a should be equal to a=1+4b

myininaya (myininaya):

\[0=2b(2-b) \\ b=0 \text{ or } 2-b=0 \]

OpenStudy (anonymous):

In the step: 0=-2b+2(-b)+1+a... 0=-4b+1+a

OpenStudy (anonymous):

not, 2b-1=a

OpenStudy (anonymous):

so a=4b-1

myininaya (myininaya):

ok you can correct it but the process is still the same we need to replace the a and c in terms of b

OpenStudy (anonymous):

alright!

myininaya (myininaya):

we can both do it let me know what you get and we can compare our answers

OpenStudy (anonymous):

would I substitute my b-values into b=-c to get c, and the 2b-2c-a-1=0 to get a?

myininaya (myininaya):

we have b=-c and you said a=4b-1

myininaya (myininaya):

so we just plug our b's into there to get our c and a

myininaya (myininaya):

did you get (a,b,c)=(3,1,-1)?

myininaya (myininaya):

my solution is backwards... my max is at (-1,-2) and my min is at (3,6) :(

myininaya (myininaya):

I will go back through it one more time to see if I could have done anything differently

OpenStudy (anonymous):

I get, a =-1, b=0, c-0 or b=3/4, a=2, and -3/4=c

myininaya (myininaya):

\[y=\frac{x^2+a}{bx+c} \\ y'=\frac{2x(bx+c)-b(x^2+a)}{(bx+c)^2} \\ y'=\frac{2bx^2+2cx-bx^2-ba}{(bx+c)^2} =\frac{bx^2+2cx-ba}{(bx+c)^2} \\ \text{we will get three equations} \\ \text{ one given by } f(-1)=-2 \text{ : } -2=\frac{1+a}{-b+c} \\ -2=\frac{1+a}{-b+c} \text{ rewrite this as : } 2b-2c=1+a \\ 2b-2c-1=a \\ \text{ now we also have } f'(3)=0 \\ \frac{9b+6c-ba}{(3b+c)^2}=0 \\ \text{ rewriting } 9b+6c-ba=0 \\ f'(-1)=0 \text{ gives } \frac{1b-2c-ba}{(-b+c)^2}=0 \\ \text{ rewriting } b-2c-ba=0 \\ \text{ so the three equations are : } \\ 2b-2c-1=a \\ 9b+6c-ba=0 \\ 1b-2c-ba=0 \\ \text{ So bottom two equations when subtracting one from the other gives } \\ 8b+8c=0 \\ b=-c \\ \text{ using } b=-c \text{ for first equation } 2b-2c-1=a \\ \text{ gives } \\ 2b+2b-1=a \text{ or } 4b-1=a \\ \text{ so let's look at the equation } 9b+6c-ba=0 \text{ and rewrite in terms of b } \\ 9b+6(-b)-b(4b-1)=0 \\ 9b-6b-4b^2+b=0 \\ 3b-4b^2+b=0 \\ 4b-4b^2 =0 \\ b=1 \text{ or } b=0 \\ \text{ so the two possibilities are : } (a,b,c)=(4(1)-1),1,-(1))=(4-1,1,-1) \\ =(3,1,-1) \\ \text{ or } (a,b,c)=(4(0)-1,0,-(0))=(-1,0,0) \\ \text{ and } (-1,0,0) \text{ makes no sense } \]

myininaya (myininaya):

http://www.wolframalpha.com/input/?i=f%28x%29%3D%28x%5E2%2B3%29%2F%28x-1%29 but this gives us what we wanted backwards if you see what I mean

OpenStudy (anonymous):

yeah i understand, why (3,1,-1) make sense cause the graph gives you the local max and min from the orginal question

OpenStudy (anonymous):

and the other option doesnt..

myininaya (myininaya):

well it give us the min and max backwards though :(

OpenStudy (anonymous):

awww

myininaya (myininaya):

your question says max at x=3 and min at x=-1 but this solution gives us max at x=-1 and min at x=3

myininaya (myininaya):

@wio what do you think here?

myininaya (myininaya):

We got (3,1,-1) but our max and min is all mixed up

myininaya (myininaya):

Kinda wonder if there is a mistake in the question Well there could be a mistake in my work. My math is sloppy today but you pretty much corrected me on any stupid mistakes.

OpenStudy (anonymous):

Hmmm

myininaya (myininaya):

I didn't do anything with the second derivative because we already had 3 equations

OpenStudy (anonymous):

So your equations are: \[ f(-1)=2\\ f'(3)=0 \land f'(-1)=0\\ f''(3)<0\land f''(-1)>0 \]

myininaya (myininaya):

with 3 unknowns

OpenStudy (anonymous):

Yeah, that is guaranteed for linear equations, but not for every type of equation.

myininaya (myininaya):

To check my solutions I used wolfram's 3 equation system solver. There seems to be no other possible solution.

OpenStudy (anonymous):

Don't worry about it... I think there was a problem with the question, cause this is only first year calc, so i don;t think it would be that difficult. I'll just ask my prof.